# Directional excitability in Hilbert spaces

### v2

@@ Abstract
We introduce a generalized excitable system in which spikes can happen in a continuum of directions, therefore drastically enriching the expressivity and control capability of the spiking dynamics. In this generalized excitable system, spiking trajectories happen in a Hilbert space with an excitable resting state at the origin and spike responses that can be triggered in any direction as a function of the system’s state and inputs. State-dependence of the spiking direction provide the system with a vanishing spiking memory trace, which enables robust tracking and integration of inputs in the spiking direction history. The model exhibits generalized forms of both Hodgkin’s Type I and Type II excitability, capturing their usual bifurcation behaviors in an abstract setting. When used as the controller of a two-dimensional navigation task, this model facilitates both the sparseness of the actuation and its sensitivity to environmental inputs. These results highlight the potential of the proposed generalized excitable model for excitable control in high- and infinite-dimensional spaces.
@@ I Introduction
In virtually all spiking models, spikes can happen in only one direction, corresponding to a binary decision: to spike or not spike. However, when looking at the collective activity of neural populations, the population spike can happen in an extremely large number of directions associated with different population spiking activity patterns that encode, e.g., sensorimotor information [9, 10] or perceptual decisions [11].
In order to increase the expressivity of simple excitable systems, this paper introduces a new, generalized model of excitable behavior, in which the spiking trajectories are occurring in a Hilbert space. The excitable resting state is located at the origin, and a spike response can be initiated in any direction of this Hilbert space. Spiking in a given direction can be considered as a fast, flexible and adaptive decision (to spike or not), over a continuum of options (all the existing directions of the Hilbert space). The direction in which a spike is generated depends on two factors: the present state of the system, as determined by the past spiking history, and the inputs to the system. The vanishing memory provided by state-dependence allows the model to filter out input noise and jittering, ensuring robust, yet reactive, response to inputs.
To ensure analytical tractability of the proposed systems dynamics, directional spiking is obtained as the feedback interconnection of a fast and a slow subsystem with two timescales:
@@
  • A slow one-dimensional dynamic, which integrates an increasing function of the norm of the state of the Hilbert-space dynamic, and uses it to modulate the potential function of the fast Hilbert-space dynamic in order to terminate its spikes and bring the Hilbert state close to the excitable state.
@@
To illustrate possible control applications of the proposed model, we test it as a directional spiking controller for a two-dimensional navigation task, in which case the spiking Hilbert space can be taken to be the two-dimensional Euclidean space. This toy example provides a simple sensorimotor testbed to explore the performance of neuromorphic controllers before their applications to more complex robotics situations. Several neuromorphic methodologies are already being developed for navigation tasks [12, 13], and the controller proposed here is akin the one considered in [14, 15]. The proposed model can be seen as a mid-point between the finite-dimensional, two-option (left or right) model used in [14] and the infinite-dimensional, infinite option (all possible planar directions) model used in [15]. The proposed model is finite-dimensional, yet capable of making spiking decision in any direction of the plane.
@@ II System definition and input-output characterization
  • The variable $x_{s}(t)\in\mathbb{R}$ of (1b) serves a modulating parameter for the shape of the potential $V$ of the gradient system (1a) (Figure 2).
@@ III-B Type II
Type II excitability exhibits a discontinuous frequency response to constant inputs $\tilde{u}$ as the input strength is increased. System (9) spikes at a finite frequency $f>f_{0}>0$ as soon as the input $\tilde{u}$ is above the rheobase $\underline{\tilde{u}}$. In this subsection, we made the Type II behavior result from a Hopf bifurcation, which makes the response of the system more resonant as compared to the saddle-node bifurcation of a Type I system. Type II excitability is better suited for synchronization and rhythmic network activity rather than fine-tuned signal encoding. Nonetheless, as the input slowly ramps up, the spiking frequency still rises, while the spikes’ width decreases, until another Hopf bifurcation occurs at a new threshold $\overline{\tilde{u}}$, beyond which the spiking stops. The following qualitative behaviors can be deduced from the phase-plane analysis of the system:
@@ IV Application to $2d$ navigation with sparse actuation
This section considers a mobile robot on a two-dimensional plane. The robot’s position is given by the vector $\mathbf{z}=(z_{x},z_{y})\in\mathbb{R}^{2}$. The robot’s acceleration, $\ddot{\mathbf{z}}$, is assumed to be controlled by a two-dimensional “all-or-none” actuator $\mbox{act}(\mathbf{u})$ that is either zero or on the unit circle of $\mathbb{R}^{2}$. Using the excitability mechanism of this paper, a controller is obtained for $\bf u$} that ensures both:
@@ IV-A Trajectory tracking with sparse actuation
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+d(\mathbf{z}^{*},\mathbf{z})}$ (14)
where $d(\cdot{},\cdot{})$ is the Euclidean distance function. In other words, the excitable model (13) receives as input the instantaneous direction of its target trajectory multiplied by the gain $k_{1}$. The regularization constant $\varepsilon>0$ appearing at the denominator eliminates the singularity at $\mathbf{z}=\mathbf{z}^{*}$. The functions $(V,g)$ are taken as in (8) and the system is parameterized by $\tau=0.01,\tau_{s}=1,\beta_{1}=3,\beta_{2}=1.5,\beta_{3}=1.5,k_{1}=0.4,\gamma=1,\alpha_{act}=3,S=0.9,\varepsilon=0.0001$. The reference trajectory is defined by $\mathbf{z}^{*}=(20+t/25)(\cos\theta_{z^{*}},\sin\theta_{z^{*}})$ with $\theta_{z^{*}}=t/100+\pi/4$, and the mobile robot and its controller are initialized with $(\mathbf{z}_{0},\dot{\mathbf{z}}_{0},\mathbf{u}_{0},u_{s,0})=0$. The trajectory $\mathbf{z}$ of the mobile robot is compared to the reference trajectory $\mathbf{z}^{*}$ in Figure 6. Both the tracking performance and the sparseness of the actuation can be appreciated in Figure 7.
@@ IV-B Responsiveness to a changing environment
In this subsection, the mobile robot’s task is to follow a reference trajectory $\mathbf{z}^{*}$ while also avoiding a set of $n_{o}$ obstacles $\{\mathbf{z}^{o_{i}}\}_{0\leq i\leq n_{o}}$, leading to the controller input
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+d(\mathbf{z}^{*},\mathbf{z})}+k_{2}\sum_{i=1}^{n_{o}}\frac{\mathbf{z}-\mathbf{z}^{o_{i}}}{d(\mathbf{z},\mathbf{z}^{o_{i}})^{2}}$ (15)
@@ V Conclusions and Perspectives
This paper introduced a novel generalization of excitable systems in a Hilbert space. The model features an excitable resting state at the origin, with spike responses triggered in any direction based on the system’s state and inputs. This system can exhibit both Hodgkin’s Type I and Type II excitability, by capturing their characteristic bifurcation behaviors in a reduced two-dimensional model. Moreover, the model’s applicability was demonstrated in a two-dimensional navigation task, where it enabled both sparse actuation and sensitivity to environmental inputs. These findings underscore the potential of this framework for neuro-inspired control schemes, particularly in high- and infinite-dimensional spaces. These more advanced applications remain to be investigated in future works.

### v3

@@ Abstract
We introduce a generalized excitable system in which spikes can happen in a continuum of directions, therefore drastically enriching the expressivity and control capability of the spiking dynamics. In this generalized excitable system, spiking trajectories happen in a Hilbert space with an excitable resting state and spike responses that can be triggered in any direction as a function of the system’s state and inputs. State-dependence of the spiking direction provide the system with a vanishing spiking memory trace, which enables robust tracking and integration of inputs in the spiking direction history. The model exhibits generalized forms of both Hodgkin’s Type I and Type II excitability, capturing their usual bifurcation behaviors in an abstract setting. In control engineering applications the proposed model facilitates both the sparseness of the actuation and its sensitivity to environmental inputs, as illustrate in a two-dimensional navigation task. These results highlight the potential of the proposed generalized excitable model for excitable control in high- and infinite-dimensional spaces.
@@ I Introduction
In virtually all spiking models, spikes can happen in only one direction, corresponding to a binary decision: to spike or not spike. However, when looking at the collective activity of neural populations, the population spike can happen in an extremely large number of directions, each associated with a different population-level spiking pattern, that encode, e.g., sensorimotor information [9, 10] or perceptual decisions [11].
In order to increase the expressivity of simple excitable systems, this paper introduces a new, generalized model of excitable behavior, in which the spiking trajectories are occurring in a Hilbert space. In the absence of inputs, the excitable resting state is located at the origin, and a spike response can be initiated in any direction of the Hilbert space in which the model is defined. Spiking in a given direction can be considered as a fast, flexible and adaptive decision (to spike or not), over a continuum of options (the directions of the Hilbert space). The direction in which a spike is generated depends on two factors: the present state of the system, as determined by the past spiking history, and the inputs to the system. The vanishing memory provided by state-dependence allows the model to filter out input noise and jittering, ensuring robust, yet reactive, excitable response to inputs.
To ensure analytical tractability of the proposed systems dynamics, directional spiking is obtained as the feedback interconnection of a fast and a slow subsystem:
@@
  • A slow one-dimensional dynamic, which integrates an increasing function of the norm of the Hilbert space state, and uses it to modulate the potential function of the Hilbert-space dynamic in order to terminate spikes and bring the Hilbert space state back to the excitable state.
@@
To illustrate possible control applications of the proposed model, we test it as a directional spiking controller for a two-dimensional navigation task, in which case the spiking Hilbert space is the two-dimensional Euclidean space. This toy example provides a simple sensorimotor testbed to explore the performance of neuromorphic controllers before their applications to more complex robotics situations. Several neuromorphic methodologies are already being developed for navigation tasks [12, 13], and the controller proposed here is akin the one considered in [14, 15]. The proposed model can be seen as a mid-point between the finite-dimensional, two-option (left or right) model used in [14] and the infinite-dimensional, infinite option (all possible planar directions) model used in [15]. The proposed model is finite-dimensional, yet capable of making spiking decision in any direction of the plane.
@@ II System definition and input-output characterization
  • The variable $x_{s}(t)\in\mathbb{R}$ of (1b) serves as modulating parameter for the shape of the potential $V$ of the gradient system (1a) (Figure 2).
@@ III-B Type II
Type II excitability exhibits a discontinuous frequency response to constant inputs $\tilde{u}$ as the input strength is increased. System (9) spikes at a finite frequency $f\geq f_{0}>0$ as soon as the input $\tilde{u}$ is above the rheobase $\underline{\tilde{u}}$. In this subsection, we made the Type II behavior result from a Hopf bifurcation, which makes the response of the system more resonant as compared to the saddle-node bifurcation of a Type I system. Type II excitability is better suited for synchronization and rhythmic network activity rather than fine-tuned signal encoding. Nonetheless, as the input slowly ramps up, the spiking frequency still rises, while the spikes’ width decreases, until another Hopf bifurcation occurs at a new threshold $\overline{\tilde{u}}$, beyond which the spiking stops. The following qualitative behaviors can be deduced from the phase-plane analysis of the system:
@@ IV Application to $2d$ navigation with sparse actuation
This section considers a mobile robot on a two-dimensional plane. The robot’s position is given by the vector $\mathbf{z}=(z_{x},z_{y})\in\mathbb{R}^{2}$. The robot’s acceleration, $\ddot{\mathbf{z}}$, is assumed to be controlled by a two-dimensional “all-or-none” actuator $\mbox{act}(\mathbf{u})\in\mathbb{R}^{2}$ that is either zero or on the unit circle of $\mathbb{R}^{2}$. Using the excitability mechanism of this paper, a controller is obtained for $\bf u$} that ensures both:
@@ IV-A Trajectory tracking with sparse actuation
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+\|\mathbf{z}^{*}-\mathbf{z}\|}$ (14)
In other words, the excitable model (13) receives as input the instantaneous direction of its target trajectory multiplied by the gain $k_{1}$. The regularization constant $\varepsilon>0$ appearing at the denominator eliminates the singularity at $\mathbf{z}=\mathbf{z}^{*}$. The functions $(V,g)$ are taken as in (8) and the system is parameterized by $\tau=0.01,\tau_{s}=1,\beta_{1}=3,\beta_{2}=1.5,\beta_{3}=1.5,k_{1}=0.4,\gamma=1,\alpha_{act}=3,S=0.9,\varepsilon=0.1$. The reference trajectory is defined by $\mathbf{z}^{*}(t)=(20+t/25)(\cos\theta^{*}(t),\sin\theta^{*}(t))$ with $\theta^{*}(t)=t/100+\pi/4$, and the mobile robot and its controller are initialized with $(\mathbf{z}_{0},\dot{\mathbf{z}}_{0},\mathbf{u}_{0},u_{s,0})=0$. The trajectory $\mathbf{z}(t)$ of the mobile robot is compared to the reference trajectory $\mathbf{z}^{*}(t)$ in Figure 6. Both the tracking performance and the sparseness of the actuation can be appreciated in Figure 7.
@@ IV-B Responsiveness to a changing environment
When the mobile robot’s task is to follow a reference trajectory $\mathbf{z}^{*}$ while also avoiding a set of $n_{o}$ obstacles $\{\mathbf{z}^{o_{i}}\}_{0\leq i\leq n_{o}}$, the excitable controller input can be modified as follows
$\mathbf{v}=k_{1}\frac{\mathbf{z}^{*}-\mathbf{z}}{\varepsilon+\|\mathbf{z}^{*}-\mathbf{z}\|}+k_{2}\sum_{i=1}^{n_{o}}\frac{\mathbf{z}-\mathbf{z}^{o_{i}}}{\|\mathbf{z}-\mathbf{z}^{o_{i}}\|^{2}}$ (15)
@@ V Conclusions and Perspectives
This paper introduced a novel generalization of excitable systems in a Hilbert space. The model features an excitable resting state close to the origin, with spike responses triggered in any direction based on the system’s state and inputs. This system can exhibit both Hodgkin’s Type I and Type II excitability, by capturing their characteristic bifurcation behaviors in a reduced two-dimensional model. Moreover, the model’s applicability was demonstrated in a two-dimensional navigation task, where it enabled both sparse actuation and sensitivity to environmental inputs. These findings underscore the potential of this framework for neuro-inspired control schemes, particularly in high- and infinite-dimensional spaces. These more advanced applications remain to be investigated in future works.